OpenStudy (anonymous):

3x-4/5-2-5x/4=1/2(x+1)

OpenStudy (anonymous):



OpenStudy (anonymous):

this is problem you have listed?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is like this (3x-4/5)-(2-5x/4)=1/2(x+1)

OpenStudy (anonymous):

good lord i cant use the stupid equation editor

OpenStudy (anonymous):

haha is there any way you could walk me through it?

OpenStudy (anonymous):

like would i have to multiply each side by 20?

OpenStudy (anonymous):

ya you could just get a common denominator on the left side by multiplying by 3x-4/5 by 4/4, and 2-5x/4 by 5/5

OpenStudy (anonymous):

then what do i do for the right side?

OpenStudy (anonymous):

im assuming the x+1 is in the numerator right

OpenStudy (anonymous):

no its 1/2(x+1)

OpenStudy (anonymous):

like 0.5(x+1)

OpenStudy (anonymous):

ok so anyways for the lhs we now have 12x-16-10+25x/20 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

except its 10-25x not 10+25x

OpenStudy (anonymous):

you didnt distribute the negative

OpenStudy (anonymous):

oh alright thanks, man i really suck at math

OpenStudy (anonymous):

simplified we now have 37x-26/20=1/2(x+1) right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

multiply rhs by 10/10 so we can get a common denominator

OpenStudy (anonymous):

alright i did

OpenStudy (anonymous):

now i have 10/20(x+1)

OpenStudy (anonymous):

37x-26/20=10/20(x+1)

OpenStudy (anonymous):

should have 10x+10/20 on the right hand side

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

.5(x+1) is equal to (x+1)/2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so now i have 37x-26/20=10x+10/20

OpenStudy (anonymous):

so were gonna subtract the right side and put it on the left so your equation should be (37x-26)/20 - ((10x+10)/20)=0

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

we have a common denominator and you can distribute the negative again so your equation should be (37x-26-10x-10)/20 = 0

OpenStudy (anonymous):

so 27x-16/20=0

OpenStudy (anonymous):

you didnt distribute the negative again

OpenStudy (anonymous):

27x+16/20=0?

OpenStudy (anonymous):

it should be (27x-36)/20=0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so now its a simple equation multiply both sides by 20 to cancel out the fraction then add the 36

OpenStudy (anonymous):

so 20 times 27x-36/20?

OpenStudy (anonymous):

ya but that will just cancel out the 20's so youll just be left with 27x-36=0

OpenStudy (anonymous):

ah i see ty

OpenStudy (anonymous):

so then add the 36 so youll have 27x=36, then divide both sides by 27 so we can get the x by itself

OpenStudy (anonymous):

x=4/3?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

ty